OpenStudy (anonymous):

Without using long division, find the remainder of (-4x3 - 9x + 10)/(1 - 2x)

OpenStudy (anonymous):

1-2x = 0, Solve for x = 1/2 Find f(1/2) = -4*(1/2)^3 - (9)*(1/2) + (10)*(1/2) Whatever answer you get after evaluating this it is your answer. f(1/2) = ?

OpenStudy (anonymous):

f(1/2)= -4*(1/8) - 9/2 +5 = -1/2 - 9/2 + 5 = -1/2 - 9/2 + 10/2 Answer?

OpenStudy (anonymous):

i don't get what you do after this do you divide them all at this point?

OpenStudy (anonymous):

is the answer 0?

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

thank you :)